# Thinking with Drafts: Speculative Temporal Reasoning for Efficient Long Video Understanding

### v1

@@ Abstract
Figure 1: (a) Paradigm comparison. Left: Conventional MLLMs with growing multi-modal contexts. Right: SpecTemp’s dual-model design—Target MLLM (7B) for reasoning, Draft MLLM (3B) for frame selection; (b) Efficiency comparison. Bubble size indicates memory footprint. Lower-right region is optimal; (c) Attention map from Qwen2.5-VL-7B [bai2025qwen2] on an input comprising video tokens sampled from four frames in LongVideoBench [wu2024longvideobench] and a language prompt; (d) Long-tailed distribution of attention scores.
@@ 1 Introduction
Long video understanding constitutes a fundamental component of human-like intelligence, enabling coherent perception, structured memory, and causal reasoning over extended temporal contexts. Recent advances in reinforcement learning with verifiable rewards (RLVR) [guo2025deepseek, shen2025vlm, yang2025r1, peng2025lmm, tan2025reason, liu2025visual, meng2025mm, yu2025perception] have substantially strengthened the reasoning ability of multi-modal large language models (MLLMs) [gpt4o, anthropic2024claude, bai2025qwen2]. Preliminary studies have pioneered the adaptation of RLVR to video understanding with task-specific reward formulations [feng2025video, park2025deepvideo, li2025videochat, chen2025scaling, wang2025time, tang2025video]. To better accommodate complex spatio-temporal understanding, recent studies have proposed the thinking-with-frames [yan2025videochat, he2025framethinker, fu2025love, li2025tempsamp, tang2025tspo, yao2025generative, zhang2025thinking] paradigm. As shown in Figure 1(a), instead of processing an entire video sequence in a single pass, this paradigm alternates between two complementary reasoning pathways: extracting temporal clues from long-horizon segments to identify salient or informative temporal regions, and examining densely sampled frames within those regions to refine local understanding. Through iterative refinement, video MLLMs progressively concentrate on informative segments, facilitating temporally grounded reasoning over extended video contexts.
@@
A recent research direction for accelerating MLLM decoding is speculative decoding [leviathan2023fast, chen2023accelerating, gagrani2024speculative], which leverages a lightweight draft model to speculate intermediate reasoning tokens and a stronger target model to verify the generated tokens. This paradigm allows significant acceleration while maintaining or even improving reasoning accuracy. Inspired by these insights, we propose a reinforcement learning based Speculative Temporal Reasoning (SpecTemp) framework to decouple temporal perception from reasoning through a cooperative dual-model design. As shown in Figure 1(a), SpecTemp delegates the frame selection process to a lightweight draft MLLM, while reserving the target MLLM for temporal reasoning and verification of the draft’s proposals. In terms of workflow, we first uniformly sample video frames and feed them into the target MLLM, which performs initial temporal reasoning and predicts temporal clues that require further attention. A lightweight draft model then conducts dense sampling within the predicted temporal regions to explore fine-grained temporal cues, while producing a compact set of sparse representative frames returned to the target model. The target model then verifies the proposed frames, decides whether sufficient information has been gathered, or iteratively triggers another round of draft speculation and verification until convergence. Through this dual-system design, our SpecTemp design mirrors human brain’s cortical collaborations [somervail2025two], where a fast perceptual subsystem (akin to the lemniscal pathway) rapidly explores the scene, and a slower cognitive module (i.e., the extralemniscal pathway) validates and integrates insights.
@@
  • Empirical Validation: Extensive experiments on video reasoning benchmarks demonstrate that SpecTemp achieves comparable or superior accuracy to VideoChat-R1.5 [yan2025videochat], while reducing inference latency by approximately 20%, highlighting its potential for real-time multimodal reasoning.
@@ 2.1 Long Video Understanding
Although MLLMs have made considerable advances in vision-language tasks, they still struggle with long-form videos exceeding minute-level durations, due to complex dynamics and limited context [zou2024seconds, tang2025video]. Early approaches primarily focused on token compression [li2024llama, song2024moviechat, shen2024longvu, man2025adacm, shu2025video] or keyframe selection [zhang2025q, wang2025videoitg, hu2025m], yet such methods often sacrifice fine-grained details due to the high compression ratio. Another line of work explores extending the LLM’s context window [chen2024longvila, shen2025long, ren2025vamba], though this does not fundamentally alleviate the high computational burden and processing costs associated with long videos. Agent-based methods [luo2024video, liu2025videomind, wang2025videochat, chen2025lvagent], on the other hand, typically divide long videos into sub-segments and leverage external tools to gather additional information. However, these approaches often require laborious customization and involve complex post-processing. As a result, an effective and scalable framework for long-form video understanding remains elusive.
@@ 2.2 Thinking with frames
With the recent success of reinforcement learning in enhancing the reasoning capabilities of MLLMs [guo2025deepseek, ouyang2022training, jaech2024openai], the thinking-with-frames paradigm [xie2025video, zhang2025thinking, wang2025video, he2025framethinker, fu2025love, yan2025videochat, tang2025tspo, yao2025k, cao2025ground] has attracted growing attention. This approach adaptively zooms into specific video clips and performs interleaved reasoning over textual and visual contexts in a unified manner, enabling more focused reasoning and improved temporal grounding. However, despite its effectiveness, it still suffers from efficiency issues: the sampled tokens from predicted temporal cues remain numerous, and the resulting visual representations are often highly redundant. To address these limitations, our proposed SpecTemp decouples temporal perception from video reasoning and employs a lightweight MLLM to extract the most relevant sparse frames, thereby achieving significant efficiency gains without sacrificing reasoning quality.
@@ 2.3 Speculative Decoding
Speculative decoding [leviathan2023fast, xia2024unlocking, chen2023accelerating] has demonstrated its effectiveness in accelerating large language models by employing a smaller draft model to rapidly generate multiple tokens that are then verified in parallel by the target LLM. Most existing approaches focus on token-level acceleration. Early works [cai2024medusa, li2024eagle, du2024glide] trained lightweight draft models to achieve high acceleration ratios through reduced latency. Self-speculative methods [xia2024swift, zhang2023draft] further improve efficiency by leveraging the original model’s internal information without introducing auxiliary models. With the recent emphasis on reasoning capabilities in LLMs [guo2025deepseek, ouyang2022training, jaech2024openai], several studies [yang2025speculative, pan2025specreason, wang2025accelerating, fu2025scaling] have shifted toward reasoning-level speculative acceleration, which differs fundamentally from traditional token-level speculative decoding. Speculative Thinking [yang2025speculative] follows this paradigm by allowing a small draft model to generate most of the reasoning process while selectively delegating challenging reasoning segments to a stronger target model. Inspired by this principle, we extend the reasoning-level speculative strategy to long video understanding, where the time-consuming frame zoom-in process in the original thinking-with-images paradigm is delegated to a small draft model, yielding significant speedups.
@@ 3.3 Policy Optimization
Optimization. Our SpecTemp undergoes a two-stage optimization process including cold-start supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT). For SFT, we separately build the high-level verification and fine-grained frame-level reasoning capabilities of the target and draft MLLMs. For RFT, we follow GRPO [shao2024deepseekmath] to model keyframe speculation and response generation as a unified policy learning process:
@@ 3.4 SpecTemp-80K Dataset Construction
Annotation. We leverage GPT-4o [gpt4o] to generate hierarchical annotations for dual-level supervision. For each video, GPT-4o produces scene-level captions describing overall narratives and key events, as well as frame-level captions capturing fine-grained visual details. These annotations provide synchronized supervision for both models: coarse temporal evidence spans for the target model’s reasoning and fine-grained frame evidence for the draft model’s selection.
@@ 4.1 Experimental Setup
Models Frames TempCompass [liu2024tempcompass] MVBench [li2024mvbench] MMVU(mc) [zhao2025mmvu] VSI-Bench [yang2025thinking] Latency(s)
@@
LongVA-7B [zhang2024long] 16 73.8 66.4 66.1 35.9 2.7
InternVL3-8B [zhu2025internvl3] 16 72.1 65.1 66.2 36.4 2.1
Video-R1 [feng2025video] 16 71.5 62.7 64.0 29.8 2.4
InternVL3-8B [zhu2025internvl3] 64 75.4 68.3 68.2 39.1 5.0
VideoChat-Flash [li2024videochat] 64 74.4 66.8 67.2 37.4 4.9
VideoChat-R1.5 [yan2025videochat] 64 77.3 70.6 70.1 39.4 5.8
Qwen2.5-VL-7B [bai2025qwen2] 16 72.2 64.1 65.7 34.3 2.1
@@
Models Frames Video-Holmes [cheng2025video] LongVideoBench [wu2024longvideobench] MLVU [zhou2024mlvu] Video-MME(wo sub) [fu2025video] Latency(s)
@@
LongVA-7B [zhang2024long] 16 43.2 56.8 47.5 60.1 4.1
InternVL3-8B [zhu2025internvl3] 16 40.1 56.1 49.0 61.8 3.9
Video-R1 [feng2025video] 16 39.8 54.6 45.6 57.4 4.7
InternVL3-8B [zhu2025internvl3] 64 44.0 60.2 51.4 62.2 9.1
VideoChat-Flash [li2024videochat] 64 43.5 59.7 51.5 63.0 10.7
VideoChat-R1.5 [yan2025videochat] 64 45.1 60.6 52.3 63.4 11.5
Qwen2.5-VL-7B [bai2025qwen2] 16 35.0 54.5 40.6 56.0 4.1
@@
Implementation Details. Our SpecTemp framework consists of a 7B-parameter target MLLM and a 3B-parameter draft MLLM, both initialized from Qwen2.5-VL [bai2025qwen2]. For the target model, we uniformly sample 10 frames during initialization. The draft model performs dense sampling at 1 fps within predicted temporal regions and selects 2 representative frames per iteration. We set the maximum iteration number to 3 to balance performance and efficiency.
Benchmarks. We conduct comprehensive evaluations across eight widely-used video understanding benchmarks, spanning both short-form and long-form scenarios. For short-form videos, we evaluate on: (1) TempCompass [liu2024tempcompass], which assesses fine-grained temporal perception through conflicting video pairs that differ only in specific temporal aspects; (2) MVBench [li2024mvbench], featuring 20 challenging tasks covering temporal skills from perception to cognition via a static-to-dynamic design; (3) MMVU [zhao2025mmvu], requiring expert-level reasoning with domain-specific knowledge across 27 subjects in four core disciplines; and (4) VSI-Bench [yang2025thinking], evaluating visual-spatial intelligence through tasks that require tracking spatial relationships across sequences. For long-form videos, we use: (1) Video-Holmes [cheng2025video], requiring models to actively locate and connect multiple scattered clues for complex multi-step reasoning; (2) LongVideoBench [wu2024longvideobench], featuring 6,678 questions with the referring reasoning paradigm for accurate retrieval from video-language interleaved inputs; (3) MLVU [zhou2024mlvu], encompassing nine diversified tasks across various video genres to assess both global and local understanding; and (4) Video-MME [fu2025video], distinguished by integrating multiple data modalities (video, subtitles, audio) across 6 visual domains with 30 subfields.
@@
Method Frames MMVU(mc) [zhao2025mmvu] Video-Holmes [cheng2025video] LongVideoBench [wu2024longvideobench] Video-MME(wo sub) [fu2025video]
@@ 6.2 More Experimental Findings
$\boldsymbol{T}_{\max}$ Video-Holmes [cheng2025video] LongVideoBench [wu2024longvideobench] Iterations Latency (s)
@@
$\boldsymbol{R_{\text{IoU}}}$ (Target) $\boldsymbol{R_{\text{visual}}}$ (Draft) Video-Holmes [cheng2025video] LongVideoBench [wu2024longvideobench]
@@
Strategy Video-Holmes [cheng2025video] LongVideoBench [wu2024longvideobench] Latency (s)
@@
Needle-in-a-Haystack evaluation. To evaluate SpecTemp’s ability to retrieve visual information from extremely long videos, we conduct a Visual Needle-In-A-Haystack (NIAH) test inspired by [zhang2024long]. We insert single-frame visual questions at various positions into hour-long video haystacks sampled at 1 FPS. The needle images are designed to be counterfactual or counter-commonsense to ensure the model cannot rely on language priors alone.
@@ 6.4 Limitations
Figure 9: Prompt template for the target model. The target model is guided to perform temporal reasoning and either predict evidence segments for further exploration or generate final answers.
Figure 10: Prompt template for the draft model. The draft model is instructed to select the most informative and diverse frames from densely sampled temporal regions.
Figure 11: Qualitative case: 1 iteration for bowling ball color identification.
Figure 12: Qualitative case: 2 iterations for sports action sequence tracking.
Figure 13: Qualitative case: 3 iterations for temporal ordering reasoning.

### v2

# arXiv:2512.00805v2[cs.CV] 26 Feb 2026

@@ Abstract
Figure 1: (a) Paradigm comparison. Left: Conventional MLLMs with growing multi-modal contexts. Right: SpecTemp’s dual-model design—Target MLLM (7B) for reasoning, Draft MLLM (3B) for frame selection; (b) Efficiency comparison. Bubble size indicates memory footprint. Lower-right region is optimal; (c) Attention map from Qwen2.5-VL-7B [2] on an input comprising video tokens sampled from four frames in LongVideoBench [66] and a language prompt; (d) Long-tailed distribution of attention scores.
@@ 1 Introduction
Long video understanding constitutes a fundamental component of human-like intelligence, enabling coherent perception, structured memory, and causal reasoning over extended temporal contexts. Recent advances in reinforcement learning with verifiable rewards (RLVR) [25, 52, 75, 49, 58, 40, 44, 79] have substantially strengthened the reasoning ability of multi-modal large language models (MLLMs) [45, 1, 2]. Preliminary studies have pioneered the adaptation of RLVR to video understanding with task-specific reward formulations [20, 48, 33, 15, 63, 60]. To better accommodate complex spatio-temporal understanding, recent studies have proposed the thinking-with-frames [71, 26, 22, 36, 59, 76, 82] paradigm. As shown in Figure 1(a), instead of processing an entire video sequence in a single pass, this paradigm alternates between two complementary reasoning pathways: extracting temporal clues from long-horizon segments to identify salient or informative temporal regions, and examining densely sampled frames within those regions to refine local understanding. Through iterative refinement, video MLLMs progressively concentrate on informative segments, facilitating temporally grounded reasoning over extended video contexts.
@@
A recent research direction for accelerating MLLM decoding is speculative decoding [29, 13, 24], which leverages a lightweight draft model to speculate intermediate reasoning tokens and a stronger target model to verify the generated tokens. This paradigm allows significant acceleration while maintaining or even improving reasoning accuracy. Inspired by these insights, we propose a reinforcement learning based Speculative Temporal Reasoning (SpecTemp) framework to decouple temporal perception from reasoning through a cooperative dual-model design. As shown in Figure 1(a), SpecTemp delegates the frame selection process to a lightweight draft MLLM, while reserving the target MLLM for temporal reasoning and verification of the draft’s proposals. In terms of workflow, we first uniformly sample video frames and feed them into the target MLLM, which performs initial temporal reasoning and predicts temporal clues that require further attention. A lightweight draft model then conducts dense sampling within the predicted temporal regions to explore fine-grained temporal cues, while producing a compact set of sparse representative frames returned to the target model. The target model then verifies the proposed frames, decides whether sufficient information has been gathered, or iteratively triggers another round of draft speculation and verification until convergence. Through this dual-system design, our SpecTemp design mirrors human brain’s cortical collaborations [56], where a fast perceptual subsystem (akin to the lemniscal pathway) rapidly explores the scene, and a slower cognitive module (i.e., the extralemniscal pathway) validates and integrates insights.
@@
  • Empirical Validation: Extensive experiments on video reasoning benchmarks demonstrate that SpecTemp achieves comparable or superior accuracy to VideoChat-R1.5 [71], while reducing inference latency by approximately 20%, highlighting its potential for real-time multimodal reasoning.
@@ 2.1 Long Video Understanding
Although MLLMs have made considerable advances in vision-language tasks, they still struggle with long-form videos exceeding minute-level durations, due to complex dynamics and limited context [90, 60]. Early approaches primarily focused on token compression [34, 57, 53, 43, 55] or keyframe selection [85, 62, 27], yet such methods often sacrifice fine-grained details due to the high compression ratio. Another line of work explores extending the LLM’s context window [16, 54, 50], though this does not fundamentally alleviate the high computational burden and processing costs associated with long videos. Agent-based methods [41, 38, 65, 12], on the other hand, typically divide long videos into sub-segments and leverage external tools to gather additional information. However, these approaches often require laborious customization and involve complex post-processing. As a result, an effective and scalable framework for long-form video understanding remains elusive.
@@ 2.2 Thinking with frames
With the recent success of reinforcement learning in enhancing the reasoning capabilities of MLLMs [25, 46, 28], the thinking-with-frames paradigm [69, 82, 61, 26, 22, 71, 59, 77, 11] has attracted growing attention. This approach adaptively zooms into specific video clips and performs interleaved reasoning over textual and visual contexts in a unified manner, enabling more focused reasoning and improved temporal grounding. However, despite its effectiveness, it still suffers from efficiency issues: the sampled tokens from predicted temporal cues remain numerous, and the resulting visual representations are often highly redundant. To address these limitations, our proposed SpecTemp decouples temporal perception from video reasoning and employs a lightweight MLLM to extract the most relevant sparse frames, thereby achieving significant efficiency gains without sacrificing reasoning quality.
@@ 2.3 Speculative Decoding
Speculative decoding [29, 68, 13] has demonstrated its effectiveness in accelerating large language models by employing a smaller draft model to rapidly generate multiple tokens that are then verified in parallel by the target LLM. Most existing approaches focus on token-level acceleration. Early works [3, 35, 19] trained lightweight draft models to achieve high acceleration ratios through reduced latency. Self-speculative methods [67, 83] further improve efficiency by leveraging the original model’s internal information without introducing auxiliary models. With the recent emphasis on reasoning capabilities in LLMs [25, 46, 28], several studies [74, 47, 64, 23] have shifted toward reasoning-level speculative acceleration, which differs fundamentally from traditional token-level speculative decoding. Speculative Thinking [74] follows this paradigm by allowing a small draft model to generate most of the reasoning process while selectively delegating challenging reasoning segments to a stronger target model. Inspired by this principle, we extend the reasoning-level speculative strategy to long video understanding, where the time-consuming frame zoom-in process in the original thinking-with-images paradigm is delegated to a small draft model, yielding significant speedups.
@@ 3.3 Policy Optimization
Optimization. Our SpecTemp undergoes a two-stage optimization process including cold-start supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT). For SFT, we separately build the high-level verification and fine-grained frame-level reasoning capabilities of the target and draft MLLMs. For RFT, we follow GRPO [51] to model keyframe speculation and response generation as a unified policy learning process:
@@ 3.4 SpecTemp-80K Dataset Construction
Annotation. We leverage GPT-4o [45] to generate hierarchical annotations for dual-level supervision. For each video, GPT-4o produces scene-level captions describing overall narratives and key events, as well as frame-level captions capturing fine-grained visual details. These annotations provide synchronized supervision for both models: coarse temporal evidence spans for the target model’s reasoning and fine-grained frame evidence for the draft model’s selection.
@@ 4.1 Experimental Setup
Models Frames TempCompass [39] MVBench [31] MMVU(mc) [86] VSI-Bench [73] Latency(s)
@@
LongVA-7B [84] 16 73.8 66.4 66.1 35.9 2.7
InternVL3-8B [89] 16 72.1 65.1 66.2 36.4 2.1
Video-R1 [20] 16 71.5 62.7 64.0 29.8 2.4
InternVL3-8B [89] 64 75.4 68.3 68.2 39.1 5.0
VideoChat-Flash [32] 64 74.4 66.8 67.2 37.4 4.9
VideoChat-R1.5 [71] 64 77.3 70.6 70.1 39.4 5.8
Qwen2.5-VL-7B [2] 16 72.2 64.1 65.7 34.3 2.1
@@
Models Frames Video-Holmes [17] LongVideoBench [66] MLVU [87] Video-MME(wo sub) [21] Latency(s)
@@
LongVA-7B [84] 16 43.2 56.8 47.5 60.1 4.1
InternVL3-8B [89] 16 40.1 56.1 49.0 61.8 3.9
Video-R1 [20] 16 39.8 54.6 45.6 57.4 4.7
InternVL3-8B [89] 64 44.0 60.2 51.4 62.2 9.1
VideoChat-Flash [32] 64 43.5 59.7 51.5 63.0 10.7
VideoChat-R1.5 [71] 64 45.1 60.6 52.3 63.4 11.5
Qwen2.5-VL-7B [2] 16 35.0 54.5 40.6 56.0 4.1
@@
Implementation Details. Our SpecTemp framework consists of a 7B-parameter target MLLM and a 3B-parameter draft MLLM, both initialized from Qwen2.5-VL [2]. For the target model, we uniformly sample 10 frames during initialization. The draft model performs dense sampling at 1 fps within predicted temporal regions and selects 2 representative frames per iteration. We set the maximum iteration number to 3 to balance performance and efficiency.
Benchmarks. We conduct comprehensive evaluations across eight widely-used video understanding benchmarks, spanning both short-form and long-form scenarios. For short-form videos, we evaluate on: (1) TempCompass [39], which assesses fine-grained temporal perception through conflicting video pairs that differ only in specific temporal aspects; (2) MVBench [31], featuring 20 challenging tasks covering temporal skills from perception to cognition via a static-to-dynamic design; (3) MMVU [86], requiring expert-level reasoning with domain-specific knowledge across 27 subjects in four core disciplines; and (4) VSI-Bench [73], evaluating visual-spatial intelligence through tasks that require tracking spatial relationships across sequences. For long-form videos, we use: (1) Video-Holmes [17], requiring models to actively locate and connect multiple scattered clues for complex multi-step reasoning; (2) LongVideoBench [66], featuring 6,678 questions with the referring reasoning paradigm for accurate retrieval from video-language interleaved inputs; (3) MLVU [87], encompassing nine diversified tasks across various video genres to assess both global and local understanding; and (4) Video-MME [21], distinguished by integrating multiple data modalities (video, subtitles, audio) across 6 visual domains with 30 subfields.
@@
Method Frames MMVU(mc) [86] Video-Holmes [17] LongVideoBench [66] Video-MME(wo sub) [21]
@@ 6.2 More Experimental Findings
$\boldsymbol{T}_{\max}$ Video-Holmes [17] LongVideoBench [66] Iterations Latency (s)
@@
$\boldsymbol{R_{\text{IoU}}}$ (Target) $\boldsymbol{R_{\text{visual}}}$ (Draft) Video-Holmes [17] LongVideoBench [66]
@@
Strategy Video-Holmes [17] LongVideoBench [66] Latency (s)
@@
Needle-in-a-Haystack evaluation. To evaluate SpecTemp’s ability to retrieve visual information from extremely long videos, we conduct a Visual Needle-In-A-Haystack (NIAH) test inspired by [84]. We insert single-frame visual questions at various positions into hour-long video haystacks sampled at 1 FPS. The needle images are designed to be counterfactual or counter-commonsense to ensure the model cannot rely on language priors alone.